OpenStudy (anonymous):

How do I solve: 6+|-10x+4|=100

OpenStudy (anonymous):

Input: 6+abs(-10 x+4) = 100 Step one: Alternate form: abs(4-10 x) = 94 Step two: Alternate form assuming x is positive: abs(2-5 x) = 47 Step three: Alternate form assuming x is real: sqrt((4-10 x)^2)+6 = 100 Final Solutions: x = 49/5

OpenStudy (anonymous):

6+|−10x+4|=100 |−10x+4|+6=100 Step 1: Add -6 to both sides. |−10x+4|+6+−6=100+−6 |−10x+4|=94 Step 2: Solve Absolute Value. |−10x+4|=94 We know either−10x+4=94or−10x+4=−94 −10x+4=94(Possibility 1) −10x+4−4=94−4(Subtract 4 from both sides) −10x=90 −10x −10 = 90 −10 (Divide both sides by -10) x=−9 −10x+4=−94(Possibility 2) −10x+4−4=−94−4(Subtract 4 from both sides) −10x=−98 −10x −10 = −98 −10 (Divide both sides by -10) x= 49 5 Answer: x=−9 or x= 49/5